# Multi-Type Preference Learning: Empowering Preference-Based Reinforcement Learning with Equal Preferences

### v1

@@ V-C ABLATION STUDY
Impact of Hyperparameters on MTPL Performance (Q3). We further analyze the impact of hyperparameters on MTPL performance within the SOTA baseline MRN [13]. Specifically, we investigate the effect of $\alpha^{Equal}$ on the tasks $Walker\_walk$ and $Pendulum\_swingup$. These tasks are selected to investigate how variations in hyperparameters influence MTPL performance under both high and low equal preference proportions. As shown in Fig. 6 a and b, our findings indicate that while learning solely from equal preferences is less effective than learning from explicit preferences, it still positively contributes to performance improvements. This supports our third research question (Q3), demonstrating that learning from equal preferences provides meaningful insights for task completion. Therefore, we set $\alpha^{explicit}$ to 1 and primarily adjust $\alpha^{Equal}$. We find that learning from both types of preferences simultaneously yields the best results, indicating that MTPL is robust to $\alpha^{Equal}$. However, we observe a slight decline in performance when $\alpha^{Equal}=1$. While we acknowledge that the convergence speed of MTPL may be slower compared to learning from any single task, we argue that achieving optimal performance with the same number of human preferences is crucial for PBRL methods. In Fig. 6 c, we explore the impact of the parameter $\alpha$ of SimTeacher on performance in the $Pendulum\_swingup$ task. In our main experiments, we set $\alpha=0.1$ and observe that as $\alpha$ increases, leading to more ambiguous judgments favoring equal preferences, our method exhibits a more pronounced improvement relative to the baseline. This finding underscores that our approach is not limited to a specific $\alpha$ value, highlighting its strong generalization capabilities.
@@ VI CONCLUSIONS
Our research introduces Multi-Type Preference Learning (MTPL) as a novel framework that enhances feedback efficiency by learning from both explicit and equal preferences. The significant performance improvements observed across various tasks indicate that MTPL effectively addresses the limitations of prior methods in utilizing equal preferences. In the future, we anticipate that MTPL will provide an innovative approach for fine-tuning large language models, facilitating better alignment with human values. This advancement has the potential to drive progress in robotics and automation, particularly in enhancing natural interactions between intelligent agents and humans.

### v2

@@ V-C ABLATION STUDIES
Impact of Hyperparameters on MTPL Performance (Q3). We further analyze the impact of hyperparameters on MTPL performance within the SOTA baseline MRN [22]. Specifically, we investigate the effect of $\alpha^{Equal}$ on the tasks $Walker\_walk$ and $Pendulum\_swingup$. These tasks are selected to investigate how variations in hyperparameters influence MTPL performance under both high and low equal preference proportions. As shown in Fig. 6 (a) and (b), our findings indicate that while learning solely from equal preferences is less effective than learning from explicit preferences, it still positively contributes to performance improvements. This supports our third research question (Q3), demonstrating that learning from equal preferences provides meaningful insights for task completion. Therefore, we set $\alpha^{explicit}$ to 1 and primarily adjust $\alpha^{Equal}$. We find that learning from both types of preferences simultaneously yields the best results, indicating that MTPL is robust to $\alpha^{Equal}$. However, we observe a slight decline in performance when $\alpha^{Equal}=1$. While we acknowledge that the convergence speed of MTPL may be slower compared to learning from any single task, we argue that achieving optimal performance with the same number of human preferences is crucial for PBRL methods. In Fig. 6 (c), we explore the impact of the parameter $\alpha$ of SimTeacher on performance in the $Pendulum\_swingup$ task. In our main experiments, we set $\alpha=0.1$ and observe that as $\alpha$ increases, leading to more ambiguous judgments favoring equal preferences, our method exhibits a more pronounced improvement relative to the baseline. This finding underscores that our approach is not limited to a specific $\alpha$ value, highlighting its strong generalization capabilities.
@@ VI CONCLUSIONS
Our research introduces Multi-Type Preference Learning (MTPL) as a novel method that enhances feedback efficiency by learning from both explicit and equal preferences. The significant performance improvements observed across various tasks indicate that MTPL effectively addresses the limitations of prior methods in utilizing equal preferences. In the future, we anticipate that MTPL will provide an innovative approach for fine-tuning large language models, facilitating better alignment with human values. This advancement has the potential to drive progress in robotics and automation, particularly in enhancing natural interactions between intelligent agents and humans.
APPENDIX
(a) Point mass easy
(b) Hopper Hop
(c) Cartpole balance sparse
(d) Ball in cup catch
(e) Reacher easy
(f) Finger turn Hard
(g) Finger turn Easy
(h) Pendulum Swingup
(i) Quadruped Walk
(j) Walker Walk
Figure 7: The learning curves of four baselines compared to MTPL across ten tasks.